OpenStudy (anonymous):

Determine the quadratic equation for the parabola: range: {y:y ≤ 9}; x-intercepts: -2 and 4

OpenStudy (saifoo.khan):

x = -2 x = 4 (x+2)(x-4) = 0 Solve.

OpenStudy (anonymous):

i have to chose from these f(x) = x2 - 2x + 8 f(x) = x2 + 2x - 8 f(x) = -x2 + 2x + 8 f(x) = -x2 - 2x + 8

OpenStudy (saifoo.khan):

You may solve this using FOIL and you will get the answer, (x+2)(x-4) = 0

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

@saifoo.khan (x+2)(x-4) = 0 is not a quadratic equation for a parabola...

OpenStudy (anonymous):

it will only give you the x intercepts of the parabola

OpenStudy (saifoo.khan):

@anonymoustwo44 , how? i never said (x+2)(x-4) = 0 is the equation. i him to solve it. (x+2)(x-4) = 0 x^2 -2x - 8 = 0 This will be it.

OpenStudy (anonymous):

x^2 -2x - 8 = 0 is not the answer. its y=x^2-2x-8

OpenStudy (saifoo.khan):

It's the same thing.. Just the way of representation is different.

OpenStudy (anonymous):

then try to graph an x^2-2x-8=0

OpenStudy (saifoo.khan):

But still, the answer you gave to the asker was still wrong.

OpenStudy (anonymous):

and so are you

OpenStudy (saifoo.khan):

Mine was better than yours. hehe.

OpenStudy (saifoo.khan):

Atleast mine dosnt have sign mistakes.

OpenStudy (anonymous):

but the fact that you set it to zero

OpenStudy (anonymous):

anyways

OpenStudy (saifoo.khan):

It's the same thing man. Just the different way of representation.

OpenStudy (saifoo.khan):

It will give us the same answer.

OpenStudy (anonymous):

ok sorry if I've been a bit harsh :D

OpenStudy (saifoo.khan):

Me too! :D